sal bought some post card. if he send the same number to each of 6 friends, he will have none left over. if he sends the same to 7 friends he will have 3 left over. how many post card did sal buy?
Answer by richwmiller(17219)

6x=(7y)+3
not enough info for a unique answer.

x = 7n+4 and y = 6n+3
Here are three possible answers
where n=1
then x=11 y=9
6*11=7*9+3
66=63+3
66 total cards
n=2
then x=18 and y=15
6*18=(7*15)+3
108=105+3
108 total cards
n=3
x=25 y=21
6*25=(7*21)+3
150=147+3
150 total cards
etc